OpenStudy (readergirl12):

The system of equations is coincident. What are the missing values? x+3y=5 4x+?y=?

OpenStudy (retireed):

What does coincident mean? Can you define it for me?

OpenStudy (readergirl12):

Coincident. Two lines or shapes that lie exactly on top of each other. Example: these two lines are coincident, only you can't see them both, because they are on top of each other source: google xD

OpenStudy (readergirl12):

@retirEEd :o

OpenStudy (mathmale):

that's right. Good. But in practical terms: You still need to find the value of he unknown coefficients. By what number must you multiply the first equation so that it (the first) starts with the same coefficient as does the second equation?

OpenStudy (readergirl12):

Um.. 5*4? Idk...

OpenStudy (anonymous):

Hello, Are you ready to help me solve it?

OpenStudy (readergirl12):

Sure.

OpenStudy (anonymous):

good. This very simple question has an assumption. Based on the assumption you can answer to it. The assumption is : x+3y=5

OpenStudy (anonymous):

so, I wanna give you another very simple question. if X=3 then how much is 2X?

OpenStudy (readergirl12):

6?

OpenStudy (anonymous):

correct! what did you do?

OpenStudy (readergirl12):

3*2

OpenStudy (anonymous):

very good. so the assumption was X=3 My question wanted the amount of 2X and you multiplied both sides of the assumption by 2. Ok? Now let's get back to your question. The assumption --> x+3y=5 Which number should be multiplied to both sides of it which makes 4X?

OpenStudy (readergirl12):

Erm..3y and 5?

OpenStudy (anonymous):

if you multiply both sides of the assumption by 2 you will have this: \[2\times(X+3Y)=2\times5 --> 2X+6Y=10\] BUT here we made 2X not 4X. So which number should be multiplied to both sides to make 4X?

OpenStudy (readergirl12):

6y and10? Or 2x and 10? v-v

OpenStudy (anonymous):

No. as you see, you can multiply both sides of equation to a SINGLE NUMBER not different numbers. For example you can multiply both sides of the assumption to 1 OR 2 OR 3 OR 4. I gave you an example of multiplying both sides by 2. We saw there that if we multiply both sides by 2 it will result in 2X+6Y=10. it is not like 4X+?Y=?. because the coefficients of Xs are different. We want them both be the same-> 4

OpenStudy (readergirl12):

Oh..Whoops.

OpenStudy (anonymous):

did you get confused?

OpenStudy (readergirl12):

Sort of..But it's alright. c:

OpenStudy (anonymous):

Ok, Do you like to multiply both sides by 3? what will be the result?

OpenStudy (readergirl12):

2*(x+3y) or 2x+6y=10?

OpenStudy (anonymous):

you multiplied by 2. :) What about 3?

OpenStudy (anonymous):

DO NOT GIVE UP! JUST DO IT...

OpenStudy (readergirl12):

I'm trying..My brain went dead. Sorry. >.<

OpenStudy (mathmale):

What are the missing values? x+3y=5 4x+?y=? We want the 2 equations to be exactly the same. Multiply the first equation by 4. Type your answer here.

OpenStudy (anonymous):

ahahaha :) relax. let me help you. Multiplying by 3 gives this: 3*(x+3y)=3*5 --> 3x+9y=15 Is it similar to 4x+?y=?

OpenStudy (readergirl12):

Um..Maybe

OpenStudy (anonymous):

3x+9y=15 4x+?y=? they are not similar, because the coefficient of Xs are different. In the first one it's 3 but in the second one it's 4. THAT'S WHY they are not similar. You got it?

OpenStudy (readergirl12):

Oh.Okay.

OpenStudy (anonymous):

:) very good. YOU ARE UNDERSTANDING IT. Now what if you multiply both sides by 4?

OpenStudy (readergirl12):

12?

OpenStudy (readergirl12):

Or 60.

OpenStudy (anonymous):

Multiplying by 4 gives: 4*(x+3y)=4*5 --> 4x+12y=20 are these two similar? 4x+12y=20 4x+?y=?

OpenStudy (readergirl12):

Yes. :o

OpenStudy (anonymous):

:D yes! they are similar. BECAUSE the coefficients of Xs are similar.

OpenStudy (readergirl12):

Omg..Thank you for being patient with me.I have no idea how I got into 8th grade...If I need more help would it be fine if I tag you? :c

OpenStudy (anonymous):

c'mon! it's really easy as you see. But before you leave I want to ask you a question. ok? Ready?

OpenStudy (readergirl12):

Yes?

OpenStudy (anonymous):

What are the missing values? 2y+5x=10 ?y+10x=?

OpenStudy (readergirl12):

Nuuuuuuu.. ;o;

OpenStudy (anonymous):

ahahaha. It's easy. What is the assumption?

OpenStudy (readergirl12):

2y+5x=10? v-v

OpenStudy (anonymous):

correct. Which number should be multiplied to both sides of it to give us something similar to ?y+10x=?

OpenStudy (readergirl12):

Um..2?

OpenStudy (anonymous):

CONGRATULATION! YES........ You got it. So what will be the result?

OpenStudy (readergirl12):

14?

OpenStudy (readergirl12):

It's wrong right.. :c

OpenStudy (anonymous):

what do you mean 14? You have to multiply 2 to both sides. 2*(2y+5x)=2*10 --> 4y+10x=20 that is similar to ?y+10x=? because the coefficients of Xs are the same. both are 10.

OpenStudy (readergirl12):

Oh gosh I'm dumb...Sorrryy.. >.<

OpenStudy (anonymous):

I think you just didn't understand what the question is looking for. The question you asked was looking for two numbers which can be replaced with two question marks. ok?

OpenStudy (readergirl12):

Okay.

OpenStudy (anonymous):

That was why we were trying to make the coefficients of Xs similar. Because then the question marks will show their values.

OpenStudy (anonymous):

If you have any difficulty with understanding it, you can ask. Do Great!

OpenStudy (readergirl12):

Thank youuu. cx